OpenStudy (anonymous):

what must be multiplied to each equation in 2m-3n=2 & m+2n=4 to elminate the variable "m" using addition?

OpenStudy (anonymous):

We must multiply the 2nd by -2

OpenStudy (jhannybean):

\[\large 2m - 3n = 2\] \[\large \color{blue}{-2}(m+2n)= 4\] Therefore 2m + (-2m) = 0 . Evaluated.

OpenStudy (anonymous):

Thanks!